• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nanda

    Log entry time 00:11:47 on November 8,2000

    Entry number 51677

    keyword=Shift Summary

    16:00 No Beam, Injector troubleshooting
    Injector problems fixed, But now machine has vacuum problems.

    17:00 Setup for Q2=4.0 GeV2
    Change spectrometer angle to 28.598
    HAC display refuses to take the new angle setting
    of 29(-38.0 vern.) . displays error#8. So, readout is still at 31.755 deg.
    Actual setting is 28.598

    21:30 Accelerator starting to tune pulsed beaam into the Hall

    21:40 CW beam On at 10uA, but short lived!
    Accelerator magnet problems.

    23:00 Beam back on at 70uA
    Start run 1149
    Calorimeter setup by Olivier

    23:30 ROC 14 Hangs, no runend, reset roc14 -> ok

    23:35 start run 1150.